### v1

# Foundation Models in Electrocardiogram: A Review

Abstract
The electrocardiogram (ECG) is ubiquitous across various healthcare domains, such as cardiac arrhythmia detection and sleep monitoring, making ECG analysis critically essential. Traditional deep learning models for ECG are task-specific, with a narrow scope of functionality and limited generalization capabilities. Recently, foundation models (FMs), also known as large pre-training models, have fundamentally reshaped the scheme of model design and representation learning, enhancing the performance across a variety of downstream tasks. This success has drawn interest in the exploration of FMs to address ECG-based medical challenges concurrently. This survey provides a timely, comprehensive and up-to-date overview of FMs for large-scale ECG-FMs. First, we offer a brief background introduction to FMs. Then, we discuss the model architectures, pre-training methods, and adaptation approaches of ECG-FMs from a methodology perspective. Despite the promising opportunities of ECG-FMs, we also outline the challenges and potential future directions. Overall, this survey aims to provide researchers and practitioners with insights into the research of ECG-FMs on theoretical underpinnings, domain-specific applications, and avenues for future exploration.
1 Introduction
Foundation models (FMs), representing a new generation of AI models trained on massive and diverse datasets, have grown rapidly across a range of fields, including language understanding [47, 78], computer vision [112], and multimodal tasks [25, 120]. Notable examples include OpenAI’s GPT-4 [2], a large language model that demonstrates state-of-the-art capabilities in natural language processing tasks like text generation and summarization, and Meta’s LLaMA-series [109, 21], which pushes the boundaries of open-source large-scale models by offering high-performance models trained on diverse data. These models are designed to generalize across multiple tasks, showing the potential for flexibility, adaptability, and scalability in a wide variety of applications [34, 76].
In the medical domain, and particularly in electrocardiography (ECG), FMs have also garnered significant attention [77, 82]. Traditionally, ECG models were built with task-specific objectives, such as detecting a specific cardiovascular condition like atrial fibrillation (AF) or premature ventricular contractions (PVCs) [56, 69]. These early models required manually labeled datasets for each condition and were limited by their narrow scope.
In contrast, ECG foundation models (ECG-FMs) present a more versatile solution, offering several advantages [97]. On the one hand, they can perform multiple tasks, such as arrhythmia detection, HRV analysis, and even more complex tasks like predicting heart failure or stratifying patients based on risk profiles [118]. For example, FMs trained on large ECG datasets, like the PTB-XL [103] or PhysioNet’s MIMIC-III [44] waveform database, have been shown to generalize across multiple tasks with high accuracy. These models leverage the diversity of the data, learning features that are transferable across tasks and patient populations, making them better suited for real-world clinical applications where data variability is common. On the other hand, ECG foundation models can incorporate multi-modal data, integrating ECG signals with other physiological or clinical data such as photoplethysmography (PPG) or electronic health records (EHRs), to provide a more comprehensive picture of patient health [86, 3].
Despite the promising prospects and rapid development of ECG-FMs, a systematic analysis of ECG-FMs from a methodological standpoint has been notably absent in prior literature. To bridge this gap, this survey aims to provide a comprehensive methodological analysis of foundation models for ECG-based clinical diagnosis. This examination will center on scrutinizing their architecture design, pre-training paradigms, and adaptation techniques. Through this endeavor, we seek to illuminate an overall picture of core elements in ECG-FMs, thereby enhancing comprehension regarding the rationale behind their efficacy and the mechanisms driving their substantial potential in ECG analysis.
In contrast to previous surveys which concentrated on general-purpose FMs or FMs covering a wide scope of medical fields, this article focuses on the ECG-based diagnosis domain. We summarize the developmental roadmap of current ECG-FMs in Figure 1, to inspire advancements and deepen comprehension in the continuously evolving landscape of FMs for ECG analysis. In brief, our major contributions are summarized as follows.
  • Comprehensive and up-to-date review. We thoroughly review the existing literature on ECG-FMs, and provide a comprehensive and up-to-date review on FMs for ECG analysis, encompassing ECG-related datasets and benchmarks, along with a wide spectrum of diagnostic tasks for cardiac diseases.
  • Novel methodology-centric taxonomy. We present a novel taxonomy that delivers a thorough analysis from a methodological standpoint of FMs for ECG analysis, accompanied by a detailed depiction of the general architectures and pipeline design.
  • Current Challenges and future research opportunities. We discuss the most challenging problems in ECG-FM research from both methodological and applicative perspectives. Additionally, we explore and emphasize potential future directions for enhancing ECG analysis through the applications of foundation models, encouraging practitioners to undertake further investigations in this field.
The rest of the article is organized as follows. Section 2 describes the brief background of FMs. Section 3 presents the methodology of ECG-FMs. Section 4 discusses the existing challenges and promising directions of ECG-FMs, and Section 5 concludes the article.
Figure 1: Roadmaps of Representative ECG-FMs.
2 Background of Foundation Models
FMs are large-scale deep-learning models trained on diverse and extensive datasets in a self-supervised or semi-supervised manner [12]. FMs generally follow the pre-training and fine-tuning paradigm, which enables the model to learn a broad range of knowledge and patterns and to leverage such prior information in various downstream tasks and applications. FMs only require considerably small-scale custom data in the fine-tuning phase, showcasing remarkable adaptability and efficiency [117]. FMs have been adopted successfully across domains such as CV, NLP, and healthcare. As the data or model size massively scales up and the available data modality dramatically increases, FMs have evolved from the initial pre-trained models like BERT to contemporary large language models (LLMs), large vision models (LVMs), vision-language models (VLMs), and time-series foundation models (TSFMs).
Large Language Models.
In the sphere of NLP, LLMs have fundamentally reshaped our understanding of syntactic and semantic information extracted from input text. Autoregressive-based LLMs, exemplified by powerful models like GPT-3 [74], PaLM [15], and ChatGPT [14, 102], have redefined text generation and question-answering capabilities by predicting subsequent words in a sequence and producing contextually coherent outputs. Additionally, the emergence of the groundbreaking LLM, LLaMA [109], represents a significant advancement, striking a balance between high performance, efficiency, and accessibility through innovative encoding tokenization and memory-efficient attention techniques. On the other hand, autoencoder-based LLMs, such as BERT [18] and its remarkable variants like T5 [91], RoBERTa [71], DistilBERT [95], and ELECTRA [17], have redefined the analysis of contextual relationships within text by leveraging bidirectional context consideration. This makes them an indispensable asset for a wide range of NLP tasks, from sentiment analysis to machine translation.
Large Vision Models.
Inspired by the success of FMs in NLP, practitioners have delved into the realm of vision FMs, encompassing LVMs and VLMs. LVMs, pre-trained on extensive image datasets, excel in comprehending image content and extracting rich semantic information. With the escalating data scale and the growing complexity of deep neural networks, LVMs have evolved from conventional CNN-based models, such as VGG [100] and EfficientNet [106], to cutting-edge Transformer-based architectures, including ViT [20], Swin Transformer [72], Linformer [114], and DeiT [108]. Moreover, many LVMs employ SSL techniques to acquire generalizable visual representations during pre-training. A prime example is MAE [38], which focuses on the correlation between visible patches and masked regions, enabling it to handle intricate visual structures and capture high-level image semantics.
Vision Language Models.
In addition to LVMs, considerable research efforts have also been channeled into the development of VLMs. VLMs typically incorporate a combined encoder-decoder architecture, leveraging an LLM as the core element, alongside a visual encoder used to transform images into a format compatible with the LLM. For instance, CLIP [89] utilizes contrastive learning methods to create fused representations for images and texts, demonstrating promising capabilities in open-vocabulary image classification. Similarly, SAM [48] stands out as a VLM that delivers impressive zero-shot image segmentation performance, enabling class-agnostic segmentation based on the image and a visual prompt specifying the targeted segments. These pre-trained VLMs can then be tailored to specific tasks by presenting them with a natural language description of the task and corresponding prompts. Recent studies have showcased the promising zero-shot performance of pre-trained VLMs across various downstream vision tasks, such as image classification, object detection, and semantic segmentation [73]. Additionally, research endeavors have also explored the development of models aligning multiple paired modalities, including image-text, video-text, and video-audio [35, 30].
Time Series Foundation Models.
The increasing richness of time-series data and its wide range of applications have thrust TSFMs into the spotlight as a promising direction for time series analysis [43, 60]. TSFMs aim to harness the power of the FM paradigm to develop generalized models adept at understanding and forecasting time series data across diverse domains. By leveraging large-scale time series datasets, TSFMs hold the promise of attaining superior performance across a spectrum of time series tasks, offering a unified framework that can expedite research and application development in this field. Based on the nature of time series, TSFMs can be classified into general-purpose TSFMs dealing with standard time series and domain-specific TSFMs focusing on particular tasks such as traffic forecasting and clinical diagnosis. General-purpose TSFMs predominantly utilize single or multiple data modalities to craft robust models targeting specific time series tasks, such as Lag-Llama [94], TimeGPT-1 [28], TTMs [23], and TSMixer [22] for zero-shot forecasting, and Voice2Series [121], TS2Vec [129], and SimMTM [19] for classification. Domain-specific TSFMs have distinct focuses and architectures tailored to the requirements of the downstream tasks. In the transportation sector, TSFMs typically employ graph structures, such as GNNs, to capture the spatio-temporal interactions within dynamic traffic systems. For instance, STGCL [70] and SPGCL [59] explore the integration of contrastive learning into spatio-temporal graph forecasting, while ST-LLM [63] combines spatio-temporal information with a partially frozen LLM to enhance traffic predictions. Additionally, TSFMs in the medical domain usually place greater emphasis on capturing the historical and hierarchical characteristics of patient data. For example, COMET [115] utilizes a multi-granularity contrastive learning framework to exploit data consistencies across different levels (e.g., observation-level and patient-level) in various medical time series, showcasing its potential benefits in complex disease diagnosis.
3 Methodology Perspective
In this section, we introduce the FMs for ECG analysis from a methodology perspective, focusing on architecture design (including backbone model and Transformer structure), and pipeline (including pre-training paradigm and fine-tuning paradigm). This discussion aims to shed light on the mechanisms of the FMs driving the effectiveness and adaptability on different ECG-based tasks.
Figure 2: Architectures of ECG-FMs.
3.1 Architecture Design
As shown in Figure 2, we first delve into the architecture of ECG-FMs, including backbone models and Transformer structure, concentrating on the underlying mechanisms that determine their capabilities and the applications across different ECG-based scenarios.
3.1.1 Transformer-based Models.
The architecture of ECG-FMs has increasingly embraced the Transformer model, which was originally developed for NLP tasks [111]. The core feature of the Transformer lies in its use of the attention mechanism, enabling the model to dynamically process different parts of the input data in parallel. The attention mechanism maps a query and a set of key-value pairs to an output, which is a weighted sum of the values weighted by the similarity between the query and each key. Since the Transformer does not rely on any convolution or recurrence, positional encoding is employed to represent the relative positions of input tokens, facilitating the capture of long-range dependencies in sequences. The scalability and efficacy of the Transformer in discerning intricate data patterns have resulted in its widespread adoption outside of NLP, such as in CV, audio, medicine, and beyond.
The original Transformer model utilizes an encoder-decoder architecture with a stack of multi-head self-attention blocks. However, the most suitable transformer model for ECG analysis is still a subject of debate, with options including encoder-only, decoder-only, and encoder-decoder ECG-FMs. Each model offers distinct properties and is suitable for various downstream tasks. The encoder-only structure processes the entire input ECG sequence at once and predicts based on the understanding of the full context, making it ideal for tasks that require interpreting the entire input data, such as cardiac disease detection [122]. On the other hand, the decoder-only architecture demonstrates strong generalization capacity for large-scale ECG pre-training and is well-suited for zero-shot inference [62, 127], ECG synthesis [16], and ECG chatbot [64, 128] applications. Additionally, prominent research combines an encoder-only model and a decoder-only model to create a sophisticated encoder-decoder structure, achieving strong performance in diverse tasks [46, 110, 128], besides using a simple Transformer [29, 135] in pre-training. The variety of architectural choices underscores the potential and necessity for further exploration in this field.
The Transformer-based backbone of pre-trained ECG-FMs typically includes seven categories: ViT-series, BERT-series, CLIP-series, GPT-series, LLaMA-series, and DAll-E-series.
  • ViT-series models. ViT was originally developed to revolutionize image recognition by directly applying the Transformer architecture to sequences of image patches [20]. Unlike CNNs that rely on local feature extraction through hierarchical layers, ViT takes a unique approach by linearly embedding images into non-overlapping vectors and treating them as input tokens for the Transformer encoder. The multi-head self-attention module in ViT empowers it to capture long-range dependencies across the entire image, setting a new standard for image analysis. ViT comes in two main variants: ViT-Base and ViT-Large. ViT-Large, with its 24 layers, 1024 hidden dimensions per layer, and 16 self-attention heads, represents a leap in architectural complexity. This deeper architecture equips ViT-Large with unparalleled capacity to capture complex patterns and long-range dependencies, making it particularly effective for handling large-scale datasets. In the realm of ECG-FM, ViT has emerged as a pre-trained model, enabling the extraction of spatial and temporal cardiac features from ECG plots or time-frequency images [80, 130]. Furthermore, ViT has also been leveraged in several studies to process ECG sequences by segmenting signals into non-overlapping time series windows [41, 96, 1, 123].
  • BERT-series models. BERT [18], a pre-trained bidirectional Transformer model, leverages contextual information from unlabeled data through masked language modeling (MLM). Its pre-training stage involves two unsupervised tasks: MLM and next sentence prediction (NSP). In the MLM task, approximately 15% of the input tokens are randomly masked with a special [MASK] token, prompting the model to predict the masked part using the surrounding context. The NSP task enables the model to comprehend sentence-level coherence by determining whether the sentence label is the ”actual next”. During the fine-tuning stage, BERT encodes a concatenated text pair using a bidirectional cross-attention mechanism between two sentences for downstream tasks. Expanding on BERT’s success in NLP, numerous BERT variants have emerged in other domains, such as BEiT [8] in CV, and BioBERT [57], ClinicalBERT [5] and BioLinkBERT [126] in medicine.
    BEiT leverages a discrete variational autoencoder (dVAE) to tokenize an image into patches and discrete visual tokens, subsequently replacing randomly masked image patches with a special mask embedding. These are then fed into a ViT to learn long-range dependencies. The pre-training task of BEiT, masked image modeling (MIM), closely resembles BERT’s MLM task, aiming to predict the visual tokens of the original image based on the encoding vectors of the corrupted image. Additionally, task-specific layers are directly appended to the pre-trained encoder to fine-tune the parameters on downstream tasks.
    BioBERT represents a groundbreaking contextual model that explores the potential of BERT within the clinical and biomedical domains [57]. It involves pretraining a BERT model using a corpus of biomedical research articles sourced from PubMed abstracts and PMC full-text articles. BioBERT utilizes the WordPiece tokenization technique [119] to represent new words with frequent subwords, effectively addressing the out-of-vocabulary issue. Requiring minimal task-specific architectural modification, BioBERT surpasses BERT across various biomedical text mining tasks, including named entity recognition, relation extraction, and question answering.
    ClinicalBERT is a specialized version of the BERT model fine-tuned on clinical texts to address the unique linguistic characteristics of medical narratives such as physician notes and discharge summaries. Unlike BioBERT, it demonstrates superior performance in named entity recognition and medical natural language inference by introducing clinical notes in the pre-training stage [5]. However, it may not be as effective on de-identification tasks, where BioBERT might outperform it due to the nature of how protected health information is masked in clinical text.
    BioLinkBERT is a pre-trained LinkBERT [126] specifically tailored for the biomedical domain, leveraging the citation links within PubMed articles. In contrast to traditional clinical BERT approaches, BioLinkBERT treats the corpus as a dynamic graph of interconnected documents, utilizing citation links to enhance the learning context. Through a dual focus on MLM and document relation prediction (DRP), BioLinkBERT captures intricate dependencies and relationships across documents, leading to superior performance in tasks requiring multi-hop reasoning and comprehension of complex inter-document relations.
  • CLIP-series models. CLIP has gained prominence as a multimodal model with zero-shot capabilities, excelling in various classification benchmarks using natural language instructions, without the need for direct optimization for each specific benchmark [90]. Trained on the extensive WebImageText [133] dataset, which consists of about 400 million pairs of images and corresponding text collected from the internet, CLIP significantly reduces reliance on expensive labeled datasets while enabling broader generalization. It utilizes separate Transformer encoders for text and image inputs, with ViT serving as the visual encoder and GPT-2 Transformer as the text encoder. By concurrently training an image and a text encoder, CLIP maximizes the similarity between embeddings of paired image-text samples while minimizing similarity with other combinations, establishing robust connections between visual representations and language text. This unique design facilitates seamless zero-shot transfer to a wide range of downstream tasks. Several following works have been proposed to enhance the pre-training performance of CLIP and customize it for different domains. For instance, DeCLIP [134], SLIP [79], and CYCLIP [31] leverage additional uni/multi-modality supervision to strengthen the model’s capabilities. Additionally, FLIP [125] introduces fine-grained alignment to the model.
    Despite the effectiveness of these CLIP variants on large-scale datasets, they encounter challenges when applied to medical image-text datasets due to their smaller scale and the potential for misclassification caused by individual differences among patients with the same disease. To tackle these challenges, practitioners have developed specialized CLIP models for the medical domain. MedCLIP [116] decouples images and texts and replaces the InfoNCE loss with a semantic matching loss based on medical knowledge to address false negatives. BiomedCLIP [131] replaces the GPT-2 Transformer with a pre-trained PubMedBERT [33] as the text encoder, showcasing exceptional performance in cross-modal retrieval, zero-shot image classification, and medical visual question answering.
  • GPT-series models. GPT is an autoregressive language model that generates human-like text using the Transformer architecture, initially introduced by [88] for natural language understanding tasks. The parameters learned during unsupervised pre-training are fine-tuned for specific tasks using annotated datasets. Trained on over 40 GB of text from public websites, GPT-2 [90] has demonstrated its capability to produce coherent and contextually relevant paragraphs. However, its practical application was limited due to performance issues. To address this, a more robust model, GPT-3 [13], was developed. GPT-3 employs a similar architecture to GPT-2 but implements alternating dense and locally banded sparse attention patterns in the transformer layers. With 175 billion parameters, GPT-3 can perform tasks without extensive retraining and significantly outperforms its predecessors in zero-shot and few-shot learning scenarios. The latest iteration, GPT-4 [2], expands the input prompt to include multimodal inputs such as image and audio, addressing a limitation of GPT-3, which only accepts text inputs. In the field of ECG analysis, several studies have utilized frozen GPT models to generate potential waveform details from medical knowledge bases [128, 127]. Other practitioners have attempted to create ECG-oriented GPT models, such as ECG-GPT [46] and SignalGPT [64]. However, GPT-series models still face challenges related to transparency and interpretability, hindering their practical implementation due to the issue of hallucinations [2, 42].
  • LLaMA-series models. LLaMA represents a significant leap forward in language generation, which is trained on a vast dataset of over 1.4 trillion tokens and enhanced with key modifications to existing pre-trained models. Apart from GPT-based LLMs, LLaMA employs innovative techniques for stability and performance, including normalizing at the input of each Transformer sub-layer, replacing the ReLU activation function with the SwiGLU [98] activation function, and utilizing rotary positional embeddings (RoPE) [104] instead of the absolute positional embeddings [15]. Furthermore, LLaMA offers specialized variants such as LLaMA-2 [109] and LLaMA-3 [21], tailored for dialogue applications, surpassing open-source chat models on various benchmarks. LLaMA-2 introduces enhanced context length and grouped-query attention (GQA) for faster inference, while LLaMA-3 optimizes the RoPE base frequency and tokenizer compression rates to reduce training compute dependency. Beyond its prowess in NLP, LLaMA-based models demonstrate exceptional performance in zero-shot ECG classification and sleep apnea detection [127].
  • DAll-E-series models. DALL-E [93] is a cutting-edge language model that employs transformer-based technology to generate images from text descriptions through a two-stage training procedure. In the initial stage, a dVAE compresses initial images into $\times×32$ grids of discrete image tokens, meticulously focusing on learning short-range pixel dependencies. Subsequently, an autoregressive Transformer models the joint distribution over text and image tokens. The overall procedure can be viewed as maximizing the evidence lower bound (ELB) on the joint likelihood of the model distribution over images, captions, and tokens. The evolution of DALL-E into variant models such as DALL-E-2 [92] and DALL-E-3 [9] has unlocked the potential to produce high-resolution images based on detailed and personalized text prompts. DALL-E-2 leverages a CLIP architecture to generate an image embedding induced by the text input and feeds the embedding to a diffusion model to create an image. DALL-E-3 far surpasses its predecessors in creating contextually relevant and visually appealing text embeds by integrating GPT-4, which demonstrates a higher sensitivity to prompt nuances while requiring less user intervention to achieve the desired output. In the realm of ECG analysis, the incorporation of DALL-E as the tokenizer in HeartBEiT [110] facilitates the conversion of ECG plots into discrete tokens, thereby enhancing the overall diagnostic process.
3.1.2 Non-Transformer-based Models.
Prior to the widespread acceptance of Transformers, a variety of traditional pre-training methods utilized models such as CNNs, RNNs, MLPs, and their variants as the backbone for pre-training. Each of these models has unique strengths and is known for its effectiveness in obtaining valuable representations.
Both CNNs and MLPs are highly regarded for their ability to effectively model spatial and temporal data. CNN-based architectures have garnered significant attention in self-supervised learning for ECG representation learning, with a notable focus on the utilization of AlexNet [50], EfficientNet [106], ResNet [39] and VGG [100] as common backbones of ECG-FMs. These approaches primarily leverage 1D convolutional operations and pooling operations. 1D convolution aids in learning short-term temporal patterns, such as heartbeat mode, while pooling can reduce data dimensionality while retaining essential features like R-peak information. Examples such as JCDCL [67] and 3KG [32] both utilize a CNN encoder to extract cardiac features, claiming superior efficiency in memory usage and processing speed while delivering competitive performance. Additionally, 1D ResNet is widely used in various pre-training modes, with Adversarial masking [10, 11], METS [58], GCL [99], and SSL-KD [85] employing different 1D ResNet variants as encoders to extract channel-dependent and periodic ECG patterns. On the other hand, MLP-based models are praised for their lightweight design, offering benefits in terms of reduced computational time and cost. For instance, CPC [74] utilizes a simple MLP to encode input sequences, demonstrating high efficiency in forecasting latent representations.
RNNs and LSTMs are widely acknowledged for their proficiency in modeling long-range temporal data. They excel at capturing dependencies across time steps, making them particularly suitable for analyzing ECG signals where the local information contained within one heartbeat cycle is generally related to earlier parts. Compared to Transformer-based FMs, RNN-based architectures offer unique strengths in ECG pre-training due to their high efficiency and low complexity in handling long ECG sequences. A notable example is the use of LSTM in CPC [74] to predict the latent representations of ECG sequences a fixed number of steps into the future. This emerging trend of integrating RNN-based models into the pre-training stage presents an invaluable opportunity for advancing ECG analysis.
3.2 Pipeline
As shown in Figure 3, we review ECG-FMs from the pipeline perspective, incorporating a wide range of model pre-training and adaptation techniques.
3.2.1 Pre-training Paradigm.
Pre-training represents a crucial first stage in developing ECG-FMs, as it equips the models with the knowledge necessary to adapt swiftly to various downstream tasks and contexts. The diverse modality of pre-training data, encompassing ECG signals, plots, and text-signal/plot pairs, allows for a wide spectrum of pre-training mechanisms in building and deploying ECG-FMs. In this survey, we introduce a new perspective that focuses on learning objectives during pre-training, categorizing existing methods for ECG-FMs into contrastive-based, reconstruction-based, generative-based, and hybrid techniques.
Figure 3: Pipeline of ECG-FMs.
Contrastive-based pre-training.
We classify the contrastive-based pre-training strategy as a comprehensive modeling of ECG representations, encompassing augmentation contrast, prediction contrast, and multimodal contrast. As shown in Table 1, augmentation contrast stands as one of the most widely used contrastive frameworks. Most ECG-FMs leverage data augmentation techniques, such as jittering, scaling, and channel masking, to generate different views of an input ECG sample, and then ensure that these augmented views are represented in a similar way in the feature space while pushing representations of different data points far apart. The key lies in constructing and utilizing the self-supervision signals by generating informative positive pairs and filtering out unsuitable negative pairs during augmentation. Due to the periodic nature of the ECG and the significance of the QRS region in containing essential cardiac information, the random heartbeat splitting and joining transformation is also meaningful for ECG augmentation contrast. For instance, sCL-ST [55] splits ECG samples into many snippets and randomizes the concatenation of these snippets to produce augmented views, significantly enhancing the robustness of learned representations across different cardiac diseases. Furthermore, augmentation contrast in the frequency domain is also feasible for ECG data. SSL-KD [85] leverages fast Fourier transform (FFT) to generate a frequency view for ECG tokens.
In contrast to data augmentation, prediction contrast takes a unique approach by predicting future information based on the present context. This method aims to maximize the preservation of mutual information between the context and the target, making it highly effective for capturing long-range dependencies in ECG data, such as heart rate variability and changes in waveform shape. SSMs [75] utilizes the contrastive predictive coding framework to capture semantic information across time for detecting cardiac arrhythmia.
Multimodal contrast generally mines data consistency by considering learning aligned representations across different modalities as the pretext task. The primary goal is to establish a shared representation space where data from various modalities, such as ECG signals and clinical notes, are aligned. During pre-training, the FM is trained to learn aligned representations across various modalities. This involves maximizing the similarity of multimodal views from the same ECG sample while minimizing the similarity of views from different samples. Existing studies typically follow CLIP framework to implement multimodal contrast pre-training. For example, sEHR-ECG-Text [52] utilizes a pre-trained BERT model to encode ECG signals, structured EHRs, and unstructured EHRs, resulting in improved performance in ECG classification, zero-shot retrieval, and out-of-distribution detection. Compared to single-modality contrast pre-training methods, multimodal contrast can generate more comprehensive and versatile representations from different modality ECG data, thereby enhancing interpretability and transferability for downstream tasks.
The three branches of contrastive-based pre-training techniques share both similarities and differences. Each branch primarily leverages InfoNCE loss as the contrastive loss function to exploit consistency in ECG samples. However, they differ in terms of the pre-training encoders they utilize. In multi-view augmentation contrast pre-training, typical models such as SimCLR and BYOL are used, while prediction contrast mainly adopts CPC. Additionally, multimodal contrast employs BERT and BEiT variants as the encoders.
Reconstruction-based pre-training.
Pre-training in ECG-FM studies generally involves the use of reconstruction techniques, such as masking and autoencoder-based methods. Masking, also known as masked representation learning, entails retaining a portion of the input and training the model to reconstruct the masked portion. Typically, this process begins by randomly masking a portion of the input ECG data, often by replacing it with a mask token or a zero-value, while keeping the rest of the data visible to the FM. For instance, adversarial masking [10, 11] involves using an image-based adversarial masking model to generate mask tokens as augmentations for each given ECG sample during pre-training, effectively reducing distortions of masked ECG signals. MassMIB [124], employs a random binary masking strategy on segmented patches of both raw ECG signals and the time-frequency spectrograms. BERTs and Masked Autoencoders (MAEs), along with their variants, are common masking-based pre-training methods in existing ECG-FM studies. BERT typically masks individual tokens, such as ECG timestamps or discrete channels, and leverages bidirectional context and self-attention to predict the masked portion through a Transformer encoder, while MAE masks larger chunks of the input, such as ECG plot patches or ECG series segments, and directly reconstructs the masked region using an encoder-decoder structure. By masking parts of the ECG data, the model can effectively learn to capture both local details (e.g., individual beats) and global structures (e.g., long-term heartbeat patterns), which is essential for robust representation learning.
Autoencoder-based reconstruction is also essential for developing pre-trained ECG-FMs. Autoencoders are capable of learning a compressed representation of input ECG data by encoding it into a lower-dimensional latent space and then reconstructing the original ECG input from this compressed representation. Models like variational autoencoders (VAEs) and denoising autoencoders (DAEs) have demonstrated outstanding performance in numerous downstream tasks, including ECG synthesis and cardiac disease detection. As an example, Auto-TTE [16] harnesses a vector-quantized VAE to compress the raw ECG signals and quantize them into discrete latent representations, thereby enhancing the resilience of synthetic ECG data through the integration of compressed ECG information.
Generative-based pre-training.
In generative-based pre-training, the FM generates future tokens by conditioning each token on the sequence of previously generated tokens. When applied to ECG analysis, this method typically utilizes Transformer-based models to encode input ECG data, such as heartbeat sequences and image patches, into high-dimensional tokens. These sequential tokens with position information enable FMs to capture temporal and spatial dependencies, as well as long-range relationships across ECG embeddings, like arrhythmias spanning multiple timestamps of recording. Current research commonly employs LLMs like GPTs as pre-trained encoders to obtain robust feature representations for downstream tasks. For instance, zero-shot RAG [127] uses LLaMA2 and GPT3.5 to convert ECG signals into prompts and integrates expert domain knowledge to guide LLMs in diagnosing arrhythmia and sleep apnea. The specific capability of generative-based pre-training methods on human instruction comprehension allows ECG-FMs to adapt to question-answering tasks, essentially acting as medical AI bots. SignalGPT [64], for example, offers an interactive interface for users to obtain essential information and fine-tunes ChatGPT by providing feedback. It uses ChatGPT as a comprehensive controller, determining the processing engine according to the signal types and analyzing the description of the given signal, based on pre-training on a massive ECG corpus.
Hybrid pre-training.
There are also several studies that utilize hybrid methods, combining contrastive-based, reconstruction-based, and generative-based techniques in the pre-training process. Hybrid pre-training has been shown to significantly improve the efficiency and robustness of representation learning by integrating the advantages of these methods. For example, GCL [99] integrates denoising autoencoder and augmentation contrast, where the decoder reconstructs the augmented views of ECG samples to reduce the noise introduced by data augmentation. CQA-ESI [128] adopts an LLM pipeline to generate detailed textual descriptions of ECGs and leverages contrastive and captioning loss to pre-train ECG encoders for enhanced representations.
3.2.2 Adaptation Paradigm.
The ECG-FM adaptation paradigm customizes ECG-FM for specific tasks or datasets, thereby enhancing its performance through the utilization of acquired knowledge about cardiac data. Previous methods are classified into four branches: linear probe, fine-tuning, Low-Rank Adaptation (LoRA), and zero-shot inference.
Linear probe is the simplest and most commonly used strategy for adapting ECG-FMs to downstream tasks. This approach involves training a linear classifier, such as logistic regression (LR), support vector machine (SVM), or random forest (RF), based on the feature representations learned by the pre-trained FM. During the adaptation process, only the weights of the linear classifier are updated through back-propagation, while the remaining FM parameters are kept frozen. The performance of this linear classifier can provide valuable insights into the usefulness of the learned representations for the task at hand. Consequently, it facilitates an assessment of the informational content within the model’s learned representations for a specific task or class, without the need for further fine-tuning of the model itself.
Fine-tuning serves as a widely utilized strategy for adapting ECG-FMs, with two prevalent approaches: full fine-tuning and additive fine-tuning. Based on the way the foundation model is used on the target dataset, there are two mainstream works: full fine-tuning and additive fine-tuning. The former involves updating all parameters of the pre-trained model, while the latter selectively updates the parameters of newly introduced layers. Similar to linear probe, Most studies employ MLP or ResNet as the additive module. Full fine-tuning necessitates substantial task-specific data and computational resources, making it suitable for tasks significantly different from the initial training tasks, with ample task-specific data available. Conversely, additive fine-tuning offers a more practical option when computational resources are limited or the target task aligns closely with the pre-training tasks.
LoRA is an efficient and scalable adaptation approach aimed at reducing the computational burden and memory requirements when adapting large-scale models for various tasks [40]. By employing matrix decomposition techniques, it can efficiently update using two small-scale low-rank matrices. The majority of pre-trained model parameters remain frozen, allowing focused training solely on the additional matrix modules. Renowned for its memory efficiency and adaptability, LoRA has gained popularity in ECG analysis. For instance, Zero-shot RAG [127] employs LoRA to accelerate the adaptation process while preventing overfitting.
Zero-shot inference has emerged as a pivotal adaptation strategy in the realm of ECG-FMs, particularly for LLM-based and VLM-based ECG-FMs. It refers to the capability of the FM to excel at tasks it was not explicitly trained for, without any additional training or fine-tuning on task-specific data. Zero-shot inference typically leverages generalized knowledge to understand and execute new tasks based solely on a description or prompt, without prior examples. For example, METS [58] employs a powerful multimodal contrastive learning framework to capitalize on the similarity between ECG embedding and text embedding, enabling it to accurately compute probabilities for zero-shot ECG classification.
Table 1: Summary of foundation model studies for ECG analysis.
[…]
[10] 11footnotemark: 1Adversarial masking ResNet1d-18 N/A Reconstruction Masking Linear LP CAC;
Age regression
PhysioNet/ CinC2020 Chapman ECG TS
[107] SleepFM CNN N/A Contrastive AC LR LP Sleep stage;
Sleep apnea
Multi-modal PSG Multi-modal PSG EEG;
ECG;
Respiratory
TS
[52] sEHR-ECG-Text BERT;
CNN;
GatorTron
EO Contrastive MC LR;
MLP
LP;
A-FT
CAC;
OOD
Mayo Clinic EHRs Mayo Clinic EHRs;
PhysioNet2020;
Chapman
ECG;
EHRs
TS+Text
[64] SignalGPT ChatGPT DO Generative AR MLP A-FT CAD;
Chatbot
CODE-test CODE-test ECG;
EEG;
EMG;
EOG
TS
[6] SimCLR;
BYOL;
VICReg
ResNet-50 N/A Contrastive AC ResNet1d-50 A-FT CAC CinC2020;
Ribeiro;
Chapman
PTB-XL ECG TS
[87] Contrastive SSL CNN1d N/A Contrastive AC MLP A-FT Stress WESAD;
RML
WESAD;
RML
ECG TS
[26] CardioGPT CLIP EO Contrastive MC Linear LP CAD GE&Philips;
NSH
GE&Philips;
NSH
ECG;
ECG reports
TF+Text
[53] DCL CNN1d
+Transformer
ED Contrastive AC Linear LP Emotion;
Blood glucose level
MAHNOB-HCI;
WESAD;
CASE;
PsPM-FR;
PsPM-RRM1-2;
ASCERTAIN
DREAMER;
D1NAMO
ECG TS
[61] ETP ResNet1d-18+ClinicalBERT EO Contrastive MC Linear LP;
ZS
CAC;
ZSC
PTB-XL;
CPSC2018
PTB-XL;
CPSC2018
ECG;
ECG reports
TS+Text
[29] Stacked Transformers Transformer ED Generative AR Linear;
LR
LP;
F-FT
CAC St-Petersburg;
PTB-XL;
Georgia;
Chapman;
Ningbo
CinC2021 ECG TS
[32] 3KG CNN1d N/A Contrastive AC Linear LP CAC PhysioNet2020 PhysioNet2020 ECG TS
[101] SimCLR;
BYOL;
SWaV
xResNet1d-50 N/A Contrastive AC Linear LP CAC PTB-XL;
Chapman
PTB-XL;
Chapman;
Ribeiro
ECG TS
[105] N/A EfficientNet-B3;
ResNet-34
N/A Contrastive AC Linear LP CAD PhysioNet/ CinC2021 PhysioNet/ CinC2021 ECG TS
[66] DLC AlexNet1d;
VGGNet1d;
ResNet1d
N/A Contrastive AC Linear LP;
F-FT
CAD NFH;
SNPH
PTB-XL;
CPSC2018;
Chapman
ECG TS
[74] CPC;
SimCLR;
BYOL;
SWaV
xResNet1d-50;
LSTM+MLP
N/A Contrastive AC Linear;
MLP
LP;
F-FT
CAD CinC2020;
Ribeiro;
Chapman
PTB-XL ECG TS
[67] 22footnotemark: 2JCDCL CNN N/A Contrastive MC Linear LP CAD NFH PTB-XL;
CPSC2018
ECG TS+Plot
[41] ECG-MAE MLE-ViT EO Reconstruction Masking Linear LP;
F-FT
CAD CinC2020;
Chapman
PTB-XL ECG TS
[99] GCL ResNet1d;
VGGNet1d
N/A Hybrid AE+AC Linear LP;
F-FT
CAC;
Denoising
SNPH PTB-XL;
Chapman;
CPSC2018
ECG TS
[96] MAESSL CNN;
ViT-Large
EO Reconstruction Masking MLP F-FT LVSD N/A N/A ECG TS
[46] ECG-GPT BEiT+GPT2 EO+DO Contrastive MC DistilBERT F-FT CAD YNHHS UK-Biobank;
PTB-XL;
Ribeiro
ECG;
ECG reports
Text+Plot
[83] 33footnotemark: 3Lead-agnostic SSL Wav2Vec+CMSC ED Contrastive AC Linear LP CAC;
Patient
CPSC;
CPSC-Extra;
PTB-XL;
Georgia;
NFH;
Chapman
CPSC;
Georgia
ECG TS
[124] 44footnotemark: 4MassMIB CNN+ViT EO Reconstruction Masking Linear F-FT CAD PTB-XL;
CPSC;
Chapman
PTB-XL;
CPSC;
Chapman
ECG TF
[51] 55footnotemark: 5N/A MSDNN+MoCo N/A Contrastive AC MSDNN F-FT CAD Cardiocloud W12L-ECG Cardiocloud W12L-ECG ECG TS
[110] 66footnotemark: 6HeartBEiT Dall-E+BEiT EO+DO Reconstruction Masking MLP F-FT LVEF;
HCM;
STEMI
MSHS MSHS;
PTB-XL
ECG Plot
[55] sCL-ST xResNet1d-50 N/A Contrastive AC Linear F-FT CAC CSPC;
INCART;
G12EC;
PTB
PTB-XL ECG TS
[16] 77footnotemark: 7Auto-TTE VQ-VAE DO Hybrid Reconstruction
+AR+GAN
N/A N/A Synthesis PTB-XL;
Sejong
PTB-XL;
Sejong
ECG;
ECG reports
TS+Text
[75] 88footnotemark: 8SSMs S4 with FCE N/A Contrastive PC MLP A-FT CAC CinC PTB-XL;
Chapman
ECG TS
[127] Zero-shot RAG LLaMA2;
GPT-3.5
DO Generative AR N/A LoRA;
ZS
CAC;
ZSC;
Sleep apnea
PTB-XL+;
Apnea-ECG
PTB-XL+;
Apnea-ECG
ECG;
ECG reports
TS+Text
[11] 99footnotemark: 9Adversarial masking ResNet1d-18 N/A Reconstruction Masking Linear LR CAC;
Gender
PhysioNet/ CinC2020 Chapman ECG TS
[135] MTECG Transformer ED Reconstruction Masking Linear A-FT AC Fuwai;
PTB-XL;
PCinC
Fuwai;
PTB-XL;
PCinC
ECG Plot
[1] LST EfficientNet1d;
ResNet1d;
ViT1d
EO Contrastive AC Linear LR Age;
BMI;
Sex
AHMS AHMS ECG;
PPG
TS
[80] ST-MEM ViT-B EO Reconstruction Masking Linear LP;
F-FT
CAC;
MIC
Chapman;
Ningbo;
CODE-15
PTB-XL;
CPSC2018;
PhysioNet2017
ECG Plot
[62] MERL-CKEPE ResNet1d-18+ Med-CPT DO Hybrid AR+Masking +MC Linear LP;
ZS
CAC;
MIC;
ZSC
MIMIC-ECG PTB-XL;
CPSC2018;
Chapman;
Ningbo
ECG;
ECG reports
TS+Text
[128] CQA-ESI GPT3.5;
ConvNext-v2;
BioLinkBERT
EO+DO Hybrid AR+MC Linear LP;
F-FT
CAC;
Patient;
Chatbot
PTB-XL;
Chapman;
MIMIC-IV-ECG
PTB-XL;
ICBEB
ECG;
ECG reports
TS+Text
[54] ISL ATT-CNN-RNN N/A Contrastive AC LR LP;
F-FT
CAC Chapman;
CPSC;
PTB-XL
Chapman;
CPSC;
PTB-XL
ECG TS
[81] SimCLR CNN1d N/A Contrastive AC Linear LP Stress N/A N/A ECG TS
[113] ASTCL CNN1d
+Transformer
ED Contrastive AC Linear LP;
F-FT
CAC;
Noise
Chapman;
PTB-XL;
CODE;
CPSC2018;
CMI
Chapman;
PTB-XL;
CODE;
CPSC2018;
CMI
ECG TS
[68] MRC CNN N/A Contrastive AC Linear LP;
F-FT
CAC NFH PTB-XL;
CPSC2018;
Chapman
ECG TS
[58] METS ResNet1d-18+ ClinicalBert EO Contrastive MC Linear LP;
ZS
ZSC MIMIC-III;
PTB-XL
PTB-XL;
MIT-BIH
ECG;
ECG reports
TS+Text
[130] MaeFE ViT1d;
Sub2D-ViT
EO Reconstruction Masking Linear LP;
F-FT
CAC CPSC2018;
PTB-XL;
Ningbo
PTB-XL;
CPSC2018
ECG TS+TF
[123] CNN+ViT Res2Net+SE
+ViT
EO Reconstruction Masking Linear F-FT CAC PTB-XL;
Chapman;
CPSC2018
PTB-XL;
Chapman;
CPSC2018
ECG TS
[85] SSL-KD SE-ResNet34 N/A Contrastive AC MLP F-FT CAC PhysioNet2020 PhysioNet2020 ECG TS+TF
[122] 1010footnotemark: 10BIOT Linear
Transformer
EO Contrastive AC MLP F-FT CAD SHHS;
PREST;
Cardiology;
CHB-MIT;
IIIC Seizure;
TUAB;
TUEV;
PTB-XL;
HAR
CHB-MIT;
IIIC Seizure;
PTB-XL;
HAR
ECG;
EEG;
HAR
TS+FD
  • a
    EO Encoder-only, DO Decoder-only, ED Encoder-decoder.
  • b
    AC Augmentation Contrast, MC Multimodal Contrast, AR Autoregressive-based,PC Prediction Contrast.
  • c
    LP Linear Probe, F-FT Full Fine-tuning, A-FT Additive Fine-tuning, ZS Zero-shot.
  • d
    CAC Cardiac Arrhythmia Classification, CAD Cardiac Abnormality Detection, MIC Myocardial Infarction Classification, Stress Stress Detection, Chatbot ECG Chatbot Assistant, Patient Patient Classification, Noise Noise Detection, ZSC Zero-shot Classification, Synthesis ECG Synthesis, Sleep stage Sleep Stage Classification, Sleep apnea Sleep Apnea Detection, OOD OOD Detection, Emotion Emotion Recognition, Blood glucose level Blood Glucose Level Monitoring, Denoising ECG Denoising, LVSD LVSD Diagnosis, LVEF LVEF Diagnosis, HCM HCM Diagnosis, STEMI STEMI Diagnosis, Gender Gender Classification, Age Age Classification, BMI BMI Classification, Sex Sex Classification.
  • e
    NFH Ningbo First Hospital, SNPH Shanghai Ninth People’s Hospital.
  • f
    TS Time Series, Plot ECG Plot, TF Time-Frequency Image, FD Frequency Domain.
  • 1
    https://github.com/jessica-bo/advmask_ecg
  • 2
    https://github.com/Aiwiscal/CDCL-ECG
  • 3
    https://github.com/Jwoo5/fairseq-signals
  • 4
    https://github.com/ysxGitHub/MassMIB
  • 5
    https://github.com/SMU-MedicalVision/ECG-Classfication
  • 6
    https://github.com/akhilvaid/HeartBEiT
  • 7
    https://github.com/TClife/text_to_ecg
  • 8
    https://github.com/tmehari/ssm_ecg
  • 9
    https://github.com/jessica-bo/advmask_ecg
  • 10
    https://github.com/ycq091044/BIOT
4 Discussion
Despite the progress made in developing ECG-FMs, several critical challenges remain that hinder their widespread adoption and effective application in clinical settings:
Limited downstream tasks were tested. The predominant focus of ECG-FMs has been on classifying cardiac diseases, with the majority of studies (37 out of 43) concentrating on the detection of arrhythmias and general abnormalities. However, this has raised concerns about the thoroughness of model evaluation and the breadth of diseases being addressed. There remains a wealth of unexplored but common cardiac disease cohorts, including hyperkalemia [27], left ventricular systolic dysfunction [7], silent atrial fibrillation [84], and hypertrophic cardiomyopathy [49], which demand our attention in developing ECG-FM. Additionally, the heavy reliance on a limited number of ECG datasets for pre-training and validation presents a significant limitation. Current research heavily leans on public datasets like PTB-XL, Physionet, CPSC, Chapman, CODE, CinC, and MIMIC. However, the scarcity of certain cardiac conditions in these datasets leads to imbalanced categories, potentially introducing bias in the detection of rare abnormalities. Therefore, it is crucial to conduct additional downstream tasks using a more diverse set of ECG data to thoroughly explore the potential of FMs in the nuanced realm of ECG analysis.
Model weights are not shared. Out of 43 studies on ECG-FM, only 3 papers [10, 122, 67] have shared their saved weight files of the pre-trained model. These papers were able to do so because they utilized public datasets. Notably, the LLM community has a strong tradition of supporting and sharing their pre-trained FMs, as evidenced by the LLaMA-series, Baichuan-series, and Qwen-series models. However, the ECG research community lacks a clear guide on how to share pre-trained weights due to data privacy and ethics concerns, which hampers the development of ECG-FMs. Additionally, data isolation remains a significant challenge, with hospitals being hesitant to share their limited ECG data and collaborate to build a comprehensive ECG database. As a result, there is a lack of diversity in available ECG data for researchers and cardiologists. Furthermore, while high-tech companies typically develop and share state-of-the-art FMs, hospitals and medical institutions struggle to make full use of their real-time ECG data due to limited computing resources. Therefore, it is crucial to encourage collaboration among these institutions to drive the development of ECG-FMs and fully leverage valuable ECG data.
Data quality is not considered. LLMs typically employ stringent data quality filtering processes to ensure that they are trained on high-quality textual data. This approach enhances the models’ ability to generate accurate and coherent outputs by learning correct syntax, semantics, and contextual nuances from clean and reliable sources. However, when this methodology is directly applied to the development of ECG-FMs, it presents significant challenges. In real clinical practice, a substantial portion of ECG data is of suboptimal quality due to various factors such as patient movement, poor electrode placement, electrical interference, and physiological artifacts like muscle contractions or respiratory movements [36, 65, 4]. If ECG-FMs are trained exclusively on high-quality, artifact-free ECG recordings, they may lack the robustness to handle the noisy and diverse data commonly encountered in clinical environments. This limitation can lead to significant degradation in model performance when deployed in real-world settings, potentially resulting in inaccurate interpretations or missed detections of critical cardiac abnormalities.
The larger the better? We don’t know. The question of whether larger models inherently lead to better performance in the realm of ECG-FMs remains unanswered. Empirical studies have undeniably shown that LLMs adhere to scaling laws—principles indicating that model performance improves predictably with increases in parameters, training data size, and computational resources [37]. These scaling laws suggest that as models grow, they can capture more intricate patterns and nuances within the data, leading to enhanced performance across various tasks. However, the applicability of these scaling laws to the ECG domain is still under investigation. ECG data possess unique characteristics, such as complex temporal dynamics and significant inter-individual variability, which may influence how model size affects performance. Unlike text data, ECG signals require models to interpret subtle physiological variations that might not benefit simply from increased parameter counts. Therefore, it is uncertain whether larger ECG models will provide better diagnostic accuracy or generalization capabilities.
ECG-FM may have the same problem as LLMs. LLMs often prove quite challenging to interpret, making it difficult to grasp the processes behind representation and output generation [132]. Despite various theoretical analyses attempting to understand pre-training and the generalization capabilities, the lack of a strong theoretical foundation remains a significant hurdle. As a result, LLMs may generate responses that are factually incorrect and misleading, yet sound plausible and coherent, a phenomenon referred to as LLM hallucinations [24]. This issue extends to ECG-FMs, posing a serious risk in the healthcare domain. In the sensitive field of ECG analysis, the potential for false knowledge or hallucinations can lead to misdiagnoses and subsequent symptoms. Moreover, ECG-FMs function as black boxes, generally providing limited insights into their decision-making processes, presenting a challenge for their practical application due to a lack of transparency and failure to meet regulatory compliance requirements. While researchers are working to address this issue, further research is necessary to enhance the fairness of ECG-FMs. Developing visualization tools that can spotlight relevant ECG features could be crucial in mitigating misinformation. Furthermore, incorporating explainable AI (xAI) techniques into ECG analysis will strengthen the trust between AI models and clinicians.
When and how are ECG-FMs ready for cardiologists? While ECG-FMs have demonstrated impressive versatility and adaptability for various cardiac-related tasks, they are still in the experimental phase. Ongoing research is focused on enhancing their performance and interpretability for clinical use. Notably, existing ECG AI tools like CardioAI and IBM Watson [45] have already gained regulatory approvals and are being implemented into clinical practice with user-friendly interfaces. However, there is still progress to be made before ECG-FMs become widely accessible and seamlessly integrated into real-time workflows. On the other hand, how can ECG-FMs really help cardiologists? First, ECG-FMs can greatly aid in cardiology report generation, potentially improving the precision of reports. Furthermore, these advanced systems have the potential to support the detection of challenging or rare cardiac conditions, offering valuable insights and suggestions in cases where cardiologists may struggle to make a confident diagnosis.
5 Conclusions
Recent advancements in FMs have been revolutionizing the field of ECG analysis. In this survey, we provide a practical guide to FMs specifically tailored for various ECG-based diagnostic scenarios from a methodology-centric perspective. We categorize ECG-FMs based on key elements such as model architecture, pre-training paradigm, and adaptation technique. Additionally, we summarize the ECG-related datasets and FM benchmarks, while addressing current challenges and outlining future directions. We believe that this survey will facilitate and inspire further innovation in the application of FMs to ECG analysis.